OpenStudy (anonymous):

Part 1: Write the equation of the line that passes through the points (–1, –4) and (2, 5). Part 2: Using complete sentences, explain whether or not it matters which point is used in the final answer. Also explain why you chose the point you did. Must Show work pleaseee

OpenStudy (anonymous):

after you find the slope, it does not matter what point you use when you write the "point-slope" formula

OpenStudy (campbell_st):

find the gradient of the line \[m = {y _{2} - y _{1}}{/} {x _{2} - x _{1}}\]

OpenStudy (campbell_st):

then use either point and substitute into the point slope equation \[y -y _{1} = m(x - x _{1})\]

OpenStudy (anonymous):

(5-(-4))/(2-(-1))=9/3=3=m if you need the answer in slope intercept form then it doesn't matter which you pick. y=3x-1